Please provide step by step instructions and print the response so I can read the response.

1) Consider the reaction

a) Use the masses of the nuclides involved to determine where it is endogenic of exoergic.
b) If it is exoergic, find the amount of energy released, if it is exoergic find the threshold energy.
Answer #1

NOTE THAT IN THESE TYPE OF PROBLEMS TRY TO WRITE AS MANY DECIMAL PLACES AS POSSIBLE TO GET ACCURATE VALUES.

Mass of Nucleus, C = 13.0033548377 u

Mass of H = 1.00782503207 u

Mass of He Nucleus = 4.00260325415 u

Mass of Be = 10.01293699 u

Therefore,

energy released is

E = 13.0033548377 + 1.00782503207 - (4.00260325415 + 10.01293699)

E = 14.01117987 - 14.01554024

E = - 0.00436037415 * 931.5

As we can see that E is negative, so, reaction is endogenic.

We will find its threshold energy

E (threshold) = (total mass * mass defect / 2 * nucleus mass (C))*931.5 MeV

E (threshold) = 4.699056e-3 * 931.5

E (threshold) = 4.377 MeV